# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v1

@@ 1.4. Related Literature
Auffinger and Cable [5] analyzed Pemantle’s Min-Plus Binary Tree (or, equivalently, first passage percolation on the critical series-parallel graph, see [18, Section 3]), proving that solutions of the associated RDE converge to a $\text{Beta}(2,1)$ distribution after a diffusive rescaling provided the initial distribution is concentrated at one. This is improved to the case of an arbitrary initial distribution in Corollary 2 below.
@@
By contrast, here the RDE’s of interest are all monotone. There is a technical issue, namely, that the result of [6] does not apply directly as the evolution (11) only makes sense as an equation posed in the space of CDF’s, whereas [6] works with evolution equations posed in the space of bounded functions. In [3], this is arguably less of an issue as the equations all involve finite differences (although the relevant details are not discussed there). Here $\mathscr{L}$ behaves like a nonlocal operator with a kernel with infinite range so at first blush one would like to know that [6] can be carried out with globally nondecreasing test functions, which is not obvious. It is, in fact, possible to adapt the method to monotone semigroups in the space of CDF’s. In the hope that it may be useful in other contexts, the result is formulated in an abstract way in Appendix B below.
When $p\neq\frac{1}{2}$, the evolution equation obtained here suggests a strategy for tackling the asymptotics of the RDE. This is of interest both in the analysis of the resistance and the distance on the series-parallel graph. The latter was recently considered by Chen, Derrida, Duquesne, and Shi in [11]. They proved that when $p>\frac{1}{2}$, the logarithm of the expected value of the distance grows at a linear rate with slope $\alpha(p)$, and they proved that $\alpha(1/2+\epsilon)\approx\sqrt{\zeta(2)\epsilon}$ as $\epsilon\to 0$.
@@ 1.5. Organization of the Paper
Section 2 discusses some preliminaries used throughout the paper. The key monotonicity result is proved there. The equation describing the evolution of CDF’s is derived in Section 3. Section 4 discusses examples that fit into the class described by (9) and (8). The main results are proved in Section 5 conditional on an extension of [6], which is presented in the appendix.
@@ 2.3. Monotonicity
$\text{if}\,\,F_{1}\leq F_{2}\,\,\text{pointwise in}\,\,\mathbb{R},\quad\text{then}\,\,TF_{1}\leq TF_{2}\,\,\text{pointwise in}\,\,\mathbb{R}.$
@@
2.4. Continuity
@@ 3. Evolution of the CDF
The next result describes the difference TFFTF-F in terms of $(Y,Z,\Theta)$. Specifically, let $\mathscr{L}$ be the operator defined on $CDF(\overline{\mathbb{R}})$ by
@@
Further, if $p=\frac{1}{2}$, then $TF(\pm\infty)=F(\pm\infty)$ for any $F\in CDF(\overline{\mathbb{R}})$.
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is smooth, then
@@
$\quad+2(1-p)\int_{x}^{\infty}\mathbf{E}[F(x_{1}-g_{-}(x_{1}-x))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
Proposition 4 has two immediate consequences that will be used in the sequel. First, it makes it relatively easy to see that $T$ commutes with the addition of constants whenever such addition makes sense.
@@
Under the assumptions of Proposition 4, if $F\in CDF(\overline{\mathbb{R}})$, then $T(F+c)=TF+c$ for any constant $c\in[-F(-\infty),1-F(+\infty)]$.
The next consequence of Proposition 4, which follows from Proposition 5 and is stated precisely in Appendix B.3, concerns the contractive property of $T$, namely,
$\|TF-TG\|_{\sup}\leq\|F-G\|_{\sup}.$
This is classical for operators with the same properties as $T$ defined instead on the larger space $B(X)$ of bounded functions on some set $X$, which is precisely the setting considered in [6]. The fact that $T$ is defined only on $CDF(\overline{\mathbb{R}})$ makes the proof slightly less obvious, but contractivity holds nonetheless.
The remainder of this section is organized as follows: In Section 3.1, Proposition 3 is proved. The formula for $\mathscr{L}$ on smooth CDF’s is then proved in Section 3.2, along with Proposition 5.
@@ 3.1. Proof of Proposition 3
Thus, since $X_{1}$ and $X_{2}$ are i.i.d. with CDF $F$, one obtains, exactly as in [18, Lemma 1.1],
@@
In [18, Lemma 1.1], it is observed that the fixed points of the map $q\mapsto pq^{2}+(1-p)(1-(1-q)^{2})$ are fundamental in the study of the asymptotics of the series-parallel graph. Yet, as in the derivation above, a fixed point of this map is nothing but a zero of the polynomial (12p)q(1q)(1-2p)q(1-q), the “reaction term” appearing in the equation.
3.1.1. Stationarity at infinity if $p=\frac{1}{2}$
If $p=\frac{1}{2}$, then
$TF(\pm\infty)-F(\pm\infty)=\mathscr{L}F(\pm\infty)$
and inspection of the formula (31) shows that $\mathscr{L}F(\pm\infty)=0$ independently of $F$. ∎
@@ 3.2. Proof of Proposition 4
Suppose that $F\in CDF(\overline{\mathbb{R}})$ is smooth. To obtain the formula (32) for $\mathscr{L}F$ in this case, it suffices to establish that
@@ 3.2.1. Case: $\Theta=1$
Condition first on the event $\{\Theta=1\}$. Before going further, it is worth noting that the assumption that $f_{+}(+\infty)=0$ implies that $X_{1}>$-$\infty$ almost surely on the event $\{Z\leq x<Y,\,\,X_{1}\leq X_{2},\,\,\Theta=1\}$ for $x\in\mathbb{R}$. This is immediate from the fact that $Y=X_{2}=Z$ on $\{X_{1}=-\infty,\,\,\Theta=1,\,\,f_{+}(+\infty)=0\}$.
@@
Therefore, since $f_{+}$ is independent of $(X_{1},X_{2})$,
@@ 3.3. Proof of Proposition 5
Suppose that $p=\frac{1}{2}$ and $F\in CDF(\overline{\mathbb{R}})$ satisfies $c_{-}\leq F\leq c_{+}$ pointwise for some $c_{-},c_{+}\in[0,1]$. Fix $c\in[-c_{-},1-c_{+}]$. The goal at this stage is to show that $T(F+c)=TF+c$.
Let $X$ be an extended-real-valued random variable with CDF $F$. Consider the CDF $F_{n}$ of the random variable $X+n^{-1}Y$, where $Y$ is a mean-zero, variance-one Gaussian, independent of $X$. Then $F_{n}$ is a smooth function and a standard coupling argument shows that both $F_{n}\to F$ and $TF_{n}\to TF$ pointwise in $\overline{\mathbb{R}}$, except for points of discontinuity, as $n\to\infty$. Therefore, since the same reasoning applies to $F+c$, it suffices to consider the case when $F$ is smooth. Yet direct inspection of formula (32) shows that $\mathscr{L}(F+c)=\mathscr{L}F$ for smooth $F$.
@@ 4. Examples
This section describes a few relevant examples, namely, those that have appeared earlier in the literature, and describes some of the useful, concrete forms of $\mathscr{L}$ that can be derived.
@@ 4.1. Resistance of the series-parallel graph
It is immediate to check that $f\in\mathcal{S}$ and sg(s)=f(g(s))s-g(s)=f(g(s)) for $s\geq\log 2$ so assumption (15) of Theorem 1 is satisfied.
@@ 4.2. Distance on the series-parallel graph
Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}<\infty\}=1$ and let $q=\mathbb{P}\{D^{(0)}>0\}$. For any $n$, $\mathbb{P}\{D^{(n)}=0\}=q$. Further,
$\lim_{N\to\infty}\mathbb{P}\{-\frac{\text{sgn}(\sigma)}{2\sqrt{|\sigma|}\sqrt{N}}\log D^{(N)}\leq x\,\mid\,D^{(N)}>0\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
@@
Let $X^{(n)}=\log D^{(n)}$ so that $\{D^{(n)}>0\}=\{X^{(n)}>-\infty\}$. By Proposition 3, $\mathbb{P}\{X^{(n)}>-\infty\}$ is constant in $n$.
Let $F_{n}$ be the CDF of $X^{(n)}$ and let $G_{n}$ be the CDF conditional on $\{X^{(n)}>-\infty\}$, that is,
$G_{n}(x)=\mathbb{P}\{X^{(n)}\leq x\,\mid\,X^{(n)}>-\infty\}.$
Notice that $G_{n}=q^{-1}(F_{n}-(1-q))$. At the same time, the strengthened version of Theorem 1 from Section 5 (Theorem 2) applies: If $F_{\epsilon}(x,t)=F_{[\epsilon^{-2}t]}(\epsilon^{-1}x)$, then the theorem implies that
$F_{\epsilon}\to F\quad\text{locally uniformly in}\,\,\mathbb{R}\times(0,\infty),$
where $F$ is the unique upper semicontinuous viscosity solution of the PDE $\partial_{t}F-\sigma|\partial_{x}F|^{2}=0$ with initial datum $F(x,0)=(1-q)+q\mathbf{1}_{[0,\infty)}(x)$. After writing $G_{\epsilon}(x,t)=G_{[\epsilon^{-2}t]}(\epsilon^{-1}x)$, it follows from this that $G_{\epsilon}\to G$, where $G$ is the unique discontinuous viscosity solution of $\partial_{t}G-q\sigma|\partial_{x}G|^{2}=0$ with $G(x,0)=\mathbf{1}_{[0,\infty)}$. The proof of Corollary 1 in Section 5 gives an explicit formula for this function, from which the claimed asymptotics of $\log D^{(N)}$ follow.
The claimed value of $\sigma$ is computed in Appendix C and is consistent with [5]. ∎
It is worth emphasizing that $D^{(n)}$ is integer-valued, hence its logarithm takes values in the set $\log\mathbb{N}$ for all $n$. For the present choice of $f_{+}$ and $f_{-}$, it is not hard to show that this remains true for any solution that starts initially in $\log\mathbb{N}$. This is worth mentioning because the equation for the CDF of such a solution takes a particularly simple form, namely, if the backward-difference operator $\nabla^{-}_{\log\mathbb{N}}$ is defined for functions on $\log\mathbb{N}$ by $\nabla^{-}_{\log\mathbb{N}}F(x)=F(x)-F(x_{-})$, where $x_{-}$ denotes the largest element of $\log\mathbb{N}$ smaller than $x$, then
$F_{n+1}-F_{n}=-p(\nabla^{-}_{\log\mathbb{N}}F)^{2}+(1-2p)F_{n}(1-F_{n})\quad\text{in}\,\,\log\mathbb{N}.$
Notice that this equation looks like a discrete-time, discrete-in-space version of (23) with $a=0$.

### v2

@@ 1.6. Related Literature
Auffinger and Cable [6] analyzed Pemantle’s Min-Plus Binary Tree (or, equivalently, first passage percolation on the critical series-parallel graph, see [21, Section 3]), proving that solutions of the associated RDE converge to a $\text{Beta}(2,1)$ distribution after a diffusive rescaling provided the initial distribution is concentrated at one. This is improved to the case of an arbitrary initial distribution in Corollary 3 below, answering one of the open problems posed therein.
The proof in [6] is based on the explicit construction of sub- and supersolutions of (11), which, by monotonicity, can be used to control solutions, or what is referred to as a barrier argument in the PDE literature. The same strategy was employed by Chen, Duquesne, and Shi in [13], who derived the same $\text{Beta}(2,2)$ limit theorem as in Corollary 1 above, albeit under slightly more restrictive assumptions on $(f_{+},f_{-})$.
@@
By contrast, here the RDE’s of interest are all monotone. There is a technical issue, namely, that the result of [7] does not apply directly as the evolution (11) only makes sense as an equation posed in the space of CDF’s, whereas [7] works with evolution equations posed in the (vector) space of bounded functions. In [3], this was arguably less of an issue as the equations all involve finite differences (although the relevant details were not discussed there). Here $\mathscr{L}$ behaves like a nonlocal operator with a kernel with infinite range so at first blush one would like to know that [7] can be carried out with globally nondecreasing test functions, which is not obvious. It is, in fact, possible to adapt the method to monotone semigroups in the space of CDF’s. In the hope that it may be useful in other contexts, the result is formulated in an abstract way in Appendix B below.
When $p\neq\frac{1}{2}$, the evolution equation obtained here and the analogy with reaction-diffusion equations suggests a strategy for tackling the asymptotics of the RDE. This is of interest both in the analysis of the resistance and the distance on the series-parallel graph. The latter was recently considered by Chen, Derrida, Duquesne, and Shi in [12]. They proved that when $p>\frac{1}{2}$, the logarithm of the expected value of the distance grows at a linear rate with slope $\alpha(p)$, and, again using barrier arguments, they proved that $\alpha(1/2+\epsilon)\approx\sqrt{\zeta(2)\epsilon}$ as $\epsilon\to 0$. Their proof is motivated through the heuristic derivation of a PDE that can be shown to be closely related to (19).
@@ 1.7. Organization of the Paper
Section 2 discusses some preliminaries used throughout the paper. The key monotonicity result is proved there. The equation describing the evolution of CDF’s is derived in Section 3. Section 4 discusses examples that fit into the class described by (9) and (8). The main results are proved in Section 5 conditional on an extension of [7] and some uniqueness results for discontinuous viscosity solutions, which are presented in the appendix.
@@ 2.1. Extended-Real-Valued Random Variables
2.2. Continuity under Vague Convergence.
It is useful to note that $T$ is continuous with respect to vague convergence. Recall that $\{F_{n}\}$ converges vaguely to $F$ if $F_{n}(x)\to F(x)$ at each point of continuity $x$ of $F$.
Proposition 1.
If $\{F_{n}\}$ is a sequence in $CDF(\overline{\mathbb{R}})$ converging vaguely to some $F$, then $\{TF_{n}\}$ converges vaguely to $TF$.
Proof.
It suffices to recall that this mode of convergence for $(F_{n})_{n\in\mathbb{N}}$ is equivalent to the existence of a probability space $\mathbb{P}$ supporting a sequence $\{(X_{1,n},X_{2,n})\}$ and a vector $(X_{1},X_{2})$ such that, for each $n$, $X_{1,n}$ and $X_{2,n}$ are i.i.d. with CDF $F_{n}$; $(X_{1},X_{2})$ are i.i.d. with CDF $F$; and $(X_{1,n},X_{2,n})\to(X_{1},X_{2})$ $\mathbb{P}$-almost surely. (Note that the possibility that $X_{1}$ or $X_{2}$ are infinite poses no difficulties here.) Letting $Y_{n}=\Phi(X_{1,n},X_{2,n})$ and $Y=\Phi(X_{1},X_{2})$, where $\Phi$ is sampled independently of $\{(X_{1,n},X_{2,n})\}$ and $(X_{1},X_{2})$, the boundedness and continuity of $f_{+}$ and $f_{-}$ are enough to deduce that $Y_{n}\to Y$ almost surely as $n\to\infty$. By definition of $T$, this implies $TF_{n}\to TF$. ∎
@@ 2.4. Monotonicity
$\text{if}\,\,F\leq G\,\,\text{pointwise in}\,\,\mathbb{R},\quad\text{then}\,\,TF\leq TG\,\,\text{pointwise in}\,\,\mathbb{R}.$
@@
Using coupling arguments very similar to the one employed above, one readily deduces that the map $T=T(p,\mathbf{P})$ is nondecreasing with respect to $p$ and $f_{+}$ and nonincreasing with respect to $f_{-}$. The result is stated next for the sake of precision.
Proposition 3.
(i) For any $p_{1}\leq p_{2}$, $T(p_{1},\mathbf{P})F\leq T(p_{2},\mathbf{P})F$ for all $F\in CDF(\overline{\mathbb{R}})$.
(ii) If $\mathbf{P}$ and $\mathbf{\tilde{P}}$ are measures on $BC([0,\infty)\times BC([0,\infty)$ for which it is possible to find a coupling of the corresponding random variables $(f_{+},f_{-})$ and (f~+,f~)($\tilde$$_{+},\tilde{f}_{-}$) such that $f_{+}\leq\tilde{f}_{+}$ and $f_{-}\geq\tilde{f}_{-}$, then $T(p,\mathbf{P})F\leq T(p,\mathbf{\tilde{P}})F$ for each $F\in CDF(\overline{\mathbb{R}})$.
2.5. Absolute Continuity
@@ 3. Evolution of the CDF
The next result describes the difference TFFTF-F in terms of $(Y,Z,\Theta)$. To that end, let $\mathscr{L}F$ be the function defined by
@@
Assume that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$. If $F\in CDF(\overline{\mathbb{R}})$ is differentiable, then
@@
$\quad+2(1-p)\int_{x}^{+\infty}\mathbf{E}[F(x_{1}-g_{-}(x_{1}-x))-F(x)]F^{\prime}(x_{1})\,dx_{1}.$
Propositions 5 and 6 immediately imply two properties of $\mathscr{L}$ that will be useful in the sequel and are stated in the next proposition.
@@
Under the assumption that $\mathbf{P}\{f_{+},f_{-}\in\mathcal{S}\}=\mathbf{P}\{f_{+}(+\infty)=f_{-}(+\infty)=0\}=1$, the function $\mathscr{L}$ has the following two properties:
  • (i)
    $(\mathscr{L}F)(+\infty)=(\mathscr{L}F)(-\infty)=0$ for each $F\in CDF(\overline{\mathbb{R}})$.
  • (ii)
    Given $F,G\in CDF(\overline{\mathbb{R}})$, if the difference FGF-G is a constant function, then $\mathscr{L}F=\mathscr{L}G$.
In particular, $T$ maps constant functions to constant functions: Its action on constants is determined by the dynamical system
$Tq-q=(1-2p)q(1-q)\quad\text{for each}\,\,q\in[0,1].$
As a consequence of property (ii) and the monotonicity of $T$, the following continuity property of $T$ follows, namely,
$\|TF-TG\|_{\sup}\leq 2(1-p)\|F-G\|_{\sup}$
In particular, $T$ is a contraction when $p=1/2$. See Appendix B.3 for the proof.
@@ 3.1. Proof of Proposition 5
Thus, since $X_{1}$ and $X_{2}$ are i.i.d. with CDF $F$, one obtains, exactly as in [21, Lemma 1.1],
@@
In [21, Lemma 1.1], it is observed that the fixed points of the map $q\mapsto pq^{2}+(1-p)(1-(1-q)^{2})$ are fundamental in the study of the asymptotics of the series-parallel graph. Yet, as in the derivation above, a fixed point of this map is nothing but a zero of the polynomial (12p)q(1q)(1-2p)q(1-q), the “reaction term” appearing in the equation.
@@ 3.2. Proof of Proposition 6
Suppose that $F\in CDF(\overline{\mathbb{R}})$ is continuously differentiable. To obtain the formula (33) for $\mathscr{L}F$ in this case, it suffices to establish that
@@ 3.2.1. Case: $\Theta=1$
Condition first on the event $\{\Theta=1\}$. Before going further, it is worth noting that the assumption that $f_{+}(+\infty)=0$ implies that $X_{1}>$-$\infty$ almost surely on the event $\{Z\leq x<Y,\,\,X_{1}\leq X_{2},\,\,\Theta=1\}$ for $x\in\mathbb{R}$. This is immediate from the fact that $Y=X_{2}=Z$ on the event $\{X_{1}=-\infty,\,\,\Theta=1,\,\,f_{+}(+\infty)=0\}$.
@@
Therefore, since $f_{+}$ is independent of $(X_{1},X_{2})$ and $\mathbb{P}\{X_{1}\leq X_{2}\}=\mathbb{P}\{X_{1}<X_{2}\}$ by the continuity of $F$,
@@ 3.3. Proof of Proposition 7
Here is the proof of (i): Fix $F\in CDF(\overline{\mathbb{R}})$. Observe that, by continuity of measure,
$\lim_{x\to+\infty}\mathbb{P}\{Z\leq x<Y\,\mid\,\Theta=1\}$ $=\lim_{x\to+\infty}\left(\mathbb{P}\{Z\leq x\,\mid\,\Theta=1\}-\mathbb{P}\{Y\leq x\,\mid\,\Theta=1\}\right)$
$=\mathbb{P}\{Z=+\infty\,\mid\,\Theta=1\}-\mathbb{P}\{Y=+\infty\,\mid\,\Theta=1\}=0$
since $\{Y=+\infty\}=\{Z=+\infty\}$ by the boundedness of $f_{+}$ and $f_{-}$. By symmetry,
$\lim_{x\to-\infty}\mathbb{P}\{Y\leq x<Z\,\mid\,\Theta=0\}=0.$
Therefore, by (32), $(\mathscr{L}F)(+\infty)=(\mathscr{L}F)(-\infty)=0$ as claimed.
Next, the proof of (ii): Suppose that $F,G\in CDF(\overline{\mathbb{R}})$ differ by a constant, hence $G=F+c$ for some $c\in\mathbb{R}$. First, if $F$ is smooth, then equation (33) applied to $F$ and $F+c$ implies $\mathscr{L}(F+c)=\mathscr{L}F$. In general, there is a sequence $\{F_{n}\}$ of smooth CDF’s that converges vaguely to $F$. By Proposition 1, $\{TF_{n}\}$ and $\{T(F_{n}+c)\}$ converge vaguely to respective limits $TF$ and $T(F+c)$. At the same time, since the polynomial P(q)=q+(12p)q(1q)P(q)=q+(1-2p)q(1-q) is smooth, the functions $\{P(F_{n})\}$ converge vaguely to $\{P(F)\}$. Therefore, for a dense set of points $x\in\mathbb{R}$,
$T(F+c)(x)-P(F+c)(x)=\lim_{n\to\infty}\mathscr{L}(F_{n}+c)(x)=\lim_{n\to\infty}\mathscr{L}F_{n}(x)=TF(x)-PF(x).$
Rearranging, this implies $T(F+c)+PF=TF+P(F+c)$ on a dense set. Since right- and lefthand sides are right-continuous (being sums of such functions), this actually implies they are equal everywhere, and, thus,
$\mathscr{L}(F+c)=T(F+c)-P(F+c)=TF-PF=\mathscr{L}F.$
Finally, consider a constant CDF $F\equiv q$ for some $q\in[0,1]$. Writing $TF=\mathcal{L}F+P(F)$ as in the first part of the proof, observe that
$(TF)(+\infty)=(\mathcal{L}F)(+\infty)+P(q)=P(q),$
and, similarly, $TF(-\infty)=P(q)$. Therefore, being nondecreasing, $TF$ must be a constant function, equal to $P(q)$. Identifying $F$ with $q$, this proves that $T$ maps $q$ to the constant $Tq=P(q)$, as claimed. ∎
@@ 4. Examples
This section describes a few relevant examples and consequences of the main theorems, Theorems 1 and 2.
@@ 4.1. Resistance of the series-parallel graph
It is immediate to check that $f\in\mathcal{S}$ and sg(s)=f(g(s))s-g(s)=f(g(s)) for $s\geq\log 2$ so the stronger assumption (16) of Theorems 1 and 2 is satisfied.
@@
The approach of this paper readily implies the following characterization of the asymptotics of the resistance $R^{(n)}$ of the critical series-parallel graph, in which resistors have i.i.d. resistances with any given law on $0,+\infty]$. The result allows for the possibility that $\mathbb{P}\{R^{(0)}=0\}>0$ or $\mathbb{P}\{R^{(0)}=+\infty\}>0$, that is, any given resistor of the graph could be open or shorted.
Corollary 2.
Assume that $p=1/2$ and let $R^{(0)}$ be any random variable taking values in $0,+\infty]$. Let $R^{(n)}$ be the resistance of the series-parallel graph at stage $n$, in which each edge has i.i.d. resistances with the same law as $R^{(0)}$, and let $q=\mathbb{P}\{0<R^{(0)}<+\infty\}$. Then $\mathbb{P}\{0<R^{(n)}<+\infty\}=q$ for each $n$ and the law of $R^{(n)}$ conditional on this event has the following limit:
$\lim_{N\to\infty}\mathbb{P}\{(36a)^{-\frac{1}{3}}N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=F_{\text{Beta}(2,2)}(q^{-\frac{1}{3}}x+\frac{1}{2}),$
where $F_{\text{Beta}(2,2)}$ is the CDF of the $\text{Beta}(2,2)$ distribution and $a=2\zeta(3)$.
Recall that the $\text{Beta}(2,2)$ distribution is determined by the PDF $6y(1-y)\mathbf{1}_{[0,1]}(y)$.
Proof.
Let $X^{(n)}=\log R^{(n)}$. As in the discussion in the introduction, $\{X^{(n)}\}$ is a solution of the RDE (9) with $f_{+}(u)=f_{-}(u)=\log(1+e^{-u})$. Let $q_{-}=\mathbb{P}\{X^{(0)}=-\infty\}$ and $q_{+}=\mathbb{P}\{X^{(0)}=+\infty\}$ so that $1-q=q_{+}+q_{-}$. Notice that
$\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{3}}X^{(0)}\leq x\}=q_{-}+q\mathbf{1}_{[0,+\infty)}(x)\quad\text{for each}\,\,x\in\mathbb{R}\setminus\{0\}.$
Let $F_{N}:\mathbb{R}\times[0,+\infty)\to[0,1]$ be the rescaled CDF of $\{X^{(n)}\}$ defined as in Theorem 2. By that theorem, $F_{N}\to F$ locally uniformly in $\mathbb{R}\times(0,+\infty)$ as $N\to+\infty$, where $F$ is the bounded discontinuous viscosity solution of the PDE
$\partial_{t}F-a|\partial_{x}F|\partial_{x}^{2}F=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad F(x,0)=q_{-}+q\mathbf{1}_{[0,+\infty)}(x).$
$F$ can be readily related to the solution of the same PDE, but with initial datum $\mathbf{1}_{[0,+\infty)}$ and a different coefficient $a$: Notice that if $G$ is defined by $G(x,t)=q^{-1}(F(x,t)-q_{-})$, then $G(x,0)=\mathbf{1}_{[0,+\infty)}(x)$ and
$\partial_{t}G-q^{2}a|\partial_{x}G|\partial_{x}^{2}G=0\quad\text{in}\,\,\mathbb{R}\times(0,+\infty),\quad G(x,0)=\mathbf{1}_{[0,+\infty)}(x).$
As in the proof of Corollary 1 (see Section 5.3) below, $G$ can be explicitly computed to be
$G(x,t)=F_{\text{Beta}(2,2)}(U(x,t)),\quad\text{where}\quad U(x,t)=\frac{x}{(36qa)^{\frac{1}{3}}}+\frac{1}{2}.$
At the same time, since $\mathbb{P}\{0<R^{(N)}<+\infty\}=q$ and $\mathbb{P}\{R^{(N)}>0\}=q_{-}$ for any $N$ by Proposition 7,
$\lim_{N\to+\infty}\mathbb{P}\{N^{-\frac{1}{3}}\log R^{(N)}\leq x\,\mid\,0<R^{(N)}<+\infty\}=\lim_{N\to+\infty}\frac{1}{q}(F_{N}(x,t)-q_{-})=G(x,t).$
@@ 4.2. Distance on the series-parallel graph
Assume that $p=\frac{1}{2}$ and $\mathbb{P}\{0\leq D^{(0)}<+\infty\}=1$ and let $q=\mathbb{P}\{D^{(0)}>0\}$. Then $\mathbb{P}\{D^{(n)}>0\}=q$ for any $n$ and the law of $D^{(n)}$ conditional on this event has the following limit:
$\lim_{N\to\infty}\mathbb{P}\{N^{-\frac{1}{2}}\log D^{(N)}\leq 2x\sqrt{|\sigma|}\,\mid\,D^{(N)}>0\}=\left\{\begin{array}[]{r l}0,&\text{if}\,\,x<0,\\ \frac{x^{2}}{q},&\text{if}\,\,0\leq x\leq\sqrt{q},\\ 1,&\text{otherwise.}\end{array}\right$.
@@
The claimed value of $\sigma$ is computed in Appendix C and is consistent with [6]. The limiting behavior of $D^{(N)}$ can be derived following the same strategy as in Corollary 2 above, once again using the fact that $\mathbb{P}\{D^{(n)}>0\}=q$ for any $n$. ∎